OpenStudy (anonymous):

how many moles of oxygen atoms are contained in the following?1.67 x 10-2mol CuSO4-5 H2O

OpenStudy (julian.pitalua):

I dont understand this CuSO4-5 H2O. Could you write it in another way?

OpenStudy (somy):

\(\huge CuSO_4•5H_2O\)

OpenStudy (somy):

there is a simple way of solving it 1. Calculate molecular mass of CuSO4•5H2O 2. Calculate molecular mass of ALL oxygen atoms in CuSO4•5H2O (which are 4 oxygen atoms in CuSO4 and 5 oxygen atoms in H2O so in total its 9 oxygen atoms, thus find molecular mass if 9 oxygen atoms) 3. Do mathematical cross multiplication Molecular mass of CuSO4•5H2O - \(1.67 \times 10^{-2}moles\) molecular mass of 9 oxygen atoms - X moles so X= \(1.67 \times 10^{-2}moles\) * molecular mass of 9 oxygen atoms / Molecular mass of CuSO4•5H2 and get the answer :3

OpenStudy (julian.pitalua):

X=(0,0167mol(CuSO45H2O) x 9,00mol(O))/1,00mol(CuSO45H2O)=0,153mol

OpenStudy (somy):

oh correction at the end i forgot O in H2O sorry :>

OpenStudy (somy):

btw its not 9 mole of O its molecular mass of 9 O atoms which means 1 O atom is 16 g/mole so 9 of O atoms is 9*16= ....g/mol of O atoms this is total molecular mass of O atoms in that compound

OpenStudy (somy):

molecular mass of CuSO4•5H2O also can be found same way using periodic table and number of each atom in this compound.

OpenStudy (julian.pitalua):

let's say that you have 1 mol of the compound, that will be 6.02xE+23 molecules, hence there are 6.023xE+23 x(9) molecules of oxygen, that is 9 moles.

OpenStudy (somy):

erm im asking for atomic mass not number of atoms as in using Avogadro constant lets say O2 O2 consists of 2 O atoms bonded together when we are asked for its molecular mass (atomic mass) then atomic mass of 1 O atom is 16 g/mole but O2 has 2 O atoms so 16*2= 36 g/mole is atomic/molecular mass of this compound O2

OpenStudy (somy):

i don't exactly know how to explain this as you guys learnt it in different way so you mix this up with moles as molecules or atoms are not same as moles but i'll try look number of moles of something * Avogadro constant= number of atoms so 1mole of something * (6.02*10^23) = number of atom of that something this means that 1 mole of that something contains 6.02*10^23 ATOMS so that many atoms is in 1 MOLE so when it comes to compounds H2O for example its not like there is 2 moles of H and 1 mole of O its 2 ATOMS of H and 1 atom of O! so you cannot take it as mole! coz mole and atom are different things! mole consists if particular number of atoms but atoms don't consist of moles so don't mix up

OpenStudy (julian.pitalua):

what when you say 1 mol of water that means that there are 6.023xE+23 molecules, hence there are 6.02xE+23 atoms of oxygen because each molecules of water has one atom of oxygen, that's 1 mol of oxygen.

OpenStudy (somy):

and who said that its 1 mole of H2O? this is the thing, it is not given how many moles of H2O are there thus you don't need to focus on that you can find mole after you balance the equation from which you get H2O using formula mole=mass/ molecular mass (thats just an example)

OpenStudy (somy):

lets say i just gave you a sample of H2O its just some amount thus you don't know how many moles is that so you cannot say that you are given 1 mole if H2O

OpenStudy (somy):

of*

OpenStudy (julian.pitalua):

i did not say that's the amount that we have, it was an example. We can't say that it is 1 mol , but it works to find the proportion that will give us the answer.

OpenStudy (somy):

idk probably works but i don't like mixing it whatsoever so i just go with my way :3